A triangle has vertices at A(2,3,1) , B(-1,1,2) , C(1,-2,3).Find the acute angle which the median to side AC makes with side BC
1
Expert's answer
2020-09-28T08:07:40-0400
"\bold{r_M}=0.5(\bold{r_A+r_C})\\=0.5((2,3,1)+(1,-2,3))=(1.5,0.5,2)"

"\cos{\alpha}=\frac{BC\cdot BM}{|BC||BM|}\\\cos{\alpha}=\frac{(1+1,-2-1,3-2)\cdot (1.5+1,0.5-1,2-2)}{|(1+1,-2-1,3-2)||(1.5+1,0.5-1,2-2)|}
\\\cos{\alpha}=\frac{(2,-3,1)\cdot (2.5,-0.5,0)}{2.55\sqrt{14}}
\\\cos{\alpha}=\frac{6.5}{9.54}\\\alpha=47\degree"